1.) Suppose a uniform random variable can be used to describe the outcome of an experiment with outcomes ranging from 20 to 80. What is the probability that this experiment results in an outcome less than 30?
Answer by stanbon(75887)

Suppose a uniform random variable can be used to describe the outcome of an experiment with outcomes ranging from 20 to 80. What is the probability that this experiment results in an outcome less than 30?
---
mean = (80+20)/2 = 50
--------------------------
6*sigma = (80-20) = 60
--------------------------
sigma = 10
---
z(30) = (30-50)/10 = -2
---
P(x < 30) = P(z < -2) = 0.0228
=================
Cheers,
Stan H.
=============